FIND THE ACCELERATION OF BODY m1 ? Is the acceleration of m1 will be different from m2 , if yes why ?

Grade:

1 Answers

Aman Bansal
11 years ago

Daer Naman,

Force = (ma) = 1 x 1, = 1N.
Mass of body A = (f/a) = 1/.48, = 2.083kg.
Mass of body B = 1/.349, = 2.865kg.
(2.865 + 2.083) = 4.948kg. 
Acceleration = (f/m) = 1/4.948, = 0.2021m/sec^2.

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert